OpenStudy (anonymous):

Find the inverse of the function f(x)=3-e^(x+4) How do I determine the domain, range, and asymptotes of both the function and the inverse function?

OpenStudy (campbell_st):

the x = 3 e^(y + 4) make y the subject x - 3 = e^(y + 4) ln(x-3) = y + 4 y = ln(x - 3) - 4

hero (hero):

The domain can only be limited if x is under a square root or in a denominator. Otherwise, it is all real numbers. There are some tricks to finding the range. It involves taking the inverse of the function and switching x and y.

OpenStudy (campbell_st):

oops should be y = ln (3 - x) -4

hero (hero):

The domain and range for the function and its inverse are inverses of each other.

hero (hero):

If that makes any sense

OpenStudy (anonymous):

so will my domain and range be the same. i got -inf., 3 as my range